# references.bib

### v1

@@ 4.1 Overview
Figure 3: An illustration of the proposed TeVAE model. Blue shapes designate trainable models, orange deterministic tensors and green distribution parameters. The shape of each tensor is designated below it. During training Z is used as the value matrix, denoted by the solid arrow, whereas during inference $\boldsymbol{\mu}_{\textbf{Z}}$ is used as the value matrix, denoted by the traced arrow. Topologically, TeVAE resembles MA-VAE [correia_ma-vae_2023].
@@ 4.2 Inference Mode
Despite the generative capabilities of VAEs, TeVAE does not leverage generation for anomaly detection. Rather than sampling a latent matrix as shown in Equation 14 during inference, sampling is disabled and only $\boldsymbol{\mu}_{\textbf{Z}}$ is taken as the input for the multi-head attention mechanism, like in [von_schleinitz_vasp_2021]. Equation 14, therefore, is replaced by Equation 19 for the forward pass.
@@
This not only accelerates inference by eliminating the sampling process but is also empirically found to be a good approximation of an averaged latent matrix if it were sampled several times like in [pereira_unsupervised_2018]. The TeVAE layout during inference is shown in Figure 3, where the traced arrow designates the information flow from the encoder to the MA mechanism.
@@ 4.4 Bypass Phenomenon
[…]
@@ 4.5 Reverse-window Process
Since the model is trained to reconstruct fixed-length windows, the same applies during inference. However, to decide whether a given measurement sequence $\mathcal{S}_{n}\in\mathbb{R}^{T_{n}\times d_{\mathcal{D}}}$ is anomalous, a continuous reconstruction of the measurement is required. A trivial way to do so would be to window the input measurement $\mathcal{S}_{n}$ using a shift of $1$, input the windows into the model and chain the last time step from each output window to obtain a continuous sequence [chen_unsupervised_2020]. Considering the BiLSTM nature of the encoder and decoder, the first and last time steps of a window can only be computed given the states from one direction, making these values, in theory, less accurate, however. To overcome this, we propose averaging matching time steps in overlapping windows, which is called mean-type reverse-window method. This is done by pre-allocating an array with NaN values, filling it, and taking the mean for each time step while ignoring the NaN values, as depicted in Figure 4.
@@
This process and the general anomaly detection process are described in Algorithm 1. The input for this process is the output distribution parameters $(\boldsymbol{\mu}_{\textbf{X}},\log\boldsymbol{\sigma}_{\textbf{X}}^{2})$, so it essentially averages the distributions and hence can only be applied to the mean and variance parameters. Consider the distributions $\mathcal{N}(\mu_{x},\sigma_{x}^{2})$ and $\mathcal{N}(\mu_{y},\sigma_{y}^{2})$, both assumed to be independant and normally distributed. The sum of both distributions results in normal distribution $\mathcal{N}(\mu_{x+y},\sigma_{x+y}^{2})$, obtained as shown in Equation 20 [lemons_introduction_2002].
@@
Sequence $\text{Sequence }\mathcal{S}_{n}\in\mathbb{R}^{T_{n}\times d_{\mathcal{D}}},\text{Threshold }\tau$
Label $\text{Label }l_{n}\text{ for sequence }\mathcal{S}_{n}$
$n_{\text{windows}}\leftarrow T_{n}-w+1$triangleright Find the total number of windows in sequence
$\boldsymbol{\mu}_{\textbf{X},\text{temp}}\leftarrow\text{zeros}(n_{\text{windows}},T_{n},d_{\mathcal{D}})+\text{NaN}$triangleright Pre-allocate NaN array to assign all windows to
$\boldsymbol{\sigma}_{\textbf{X},\text{temp}}^{2}\leftarrow\text{zeros}(n_{\text{windows}},T_{n},d_{\mathcal{D}})+\text{NaN}$triangleright Pre-allocate NaN array to assign all windows to
@@
     $\boldsymbol{\mu}_{\textbf{X},\text{temp}}[i,i:i+w]\leftarrow\boldsymbol{\mu}_{\textbf{X}}$triangleright Assign mean parameter to pre-allocated array
     $\boldsymbol{\sigma}_{\textbf{X},\text{temp}}^{2}[i,i:i+w]\leftarrow\boldsymbol{\sigma}_{\textbf{X}}^{2}$triangleright Assign variance parameter to pre-allocated array
@@
$\boldsymbol{\mu}_{\mathcal{S}}\leftarrow\text{nanmean}(\boldsymbol{\mu}_{\textbf{X},\text{temp}},0)$triangleright Take nanmean along axis 0
$\boldsymbol{\sigma}_{\mathcal{S}}^{2}\leftarrow\text{nanmean}(\boldsymbol{\sigma}_{\textbf{X},\text{temp}}^{2},0)$triangleright Take nanmean along axis 0
@@
The theoretical delay $\delta_{\text{theory}}$ associated with each of the reverse-window processes can be discussed ahead of Section 5, however. $\delta_{\text{theory}}$ is defined as the intrinsic delay introduced by each reverse-windowing method. To illustrate the theoretical delay $\delta_{\text{theory}}$, it is plotted against time $t$ to demonstrate the delay for each of the reverse-window processes, shown in Figure 5. For the last-type reverse-window method during $0<t<w$, no time steps can be evaluated until a full window can be formed, streamed and evaluated, introducing a theoretical delay $\delta_{\text{theory}}=w$. This property is intrinsic to approaches based on fixed-length windows rather than variable-length sequences. At time step $t=w$, however, all time steps $0<t<w$ are evaluated and output at the same time. For $w<t<T$, however, the last time step of each window corresponds the current real-world time step, i.e. $\delta_{\text{theory}}=0$ for $w<t<T$. As mentioned above, the lack of evaluation until $t=w$ is natural to any window-based approach and hence the first-type and mean-type reverse-window methods show the same behaviour. In contrast to last-type reverse-windowing, however, these methods only output the first value of each window i.e. evaluation is $\delta_{\text{theory}}=w$ time steps behind for $0<t<T-w$. At $t=T$, though, the time steps $T-w<t<T$ are all output at the same time, meaning that $\delta_{\text{theory}}=0$ and no extra time is needed for evaluation after streaming ends. It should be noted, however, that in online time-series anomaly detection the last-type reverse-window method is in theory faster than the other two types. Consider a sub-sequence anomaly starting at time step $w+2$, designated by the red line in Figure 5. Apart from the time required for inference, the last-type reverse-window method can detect the anomaly without a theoretical delay $\delta_{\text{theory}}$, during $w<t<T$, whereas the other two methods can only detect the anomaly $\delta_{\text{theory}}=w$ time steps later. For $0<t<T-w$, a theoretical delay of $\delta_{\text{theory}}=w$ time steps is, therefore, the absolute best that first-type and mean-type reverse-windowing can achieve, however, while the best delay the first-type can achieve is $\delta_{\text{theory}}=0$, it will be higher in reality, perhaps higher than $w$ time steps.
@@ 4.6 Root-cause Analysis
The first known to the authors is proposed by [zhang_deep_2019]. The approach does not use time series but instead, the resulting correlation matrices and the anomaly score is given as the difference between the input and reconstructed correlation matrices. Hence, as a mean of providing root-cause information they rank each channel by the number of high anomaly scores for every other channel. Root-cause identification is then quantified by the recall for the top $k$ channels, in this case three, although little information on how exactly this is calculated is provided. Another issue with this metric is that is not parameter-free, since $k$ needs to be set manually. Furthermore, the method used is specific to approaches using the same type of correlation matrices, which is rare in literature.
Approaches using the negative log-likelihood as the anomaly score obtain it from a multivariate output distribution with a diagonal covariance. While the resulting anomaly score is a univariate time series, it is referred to as multivariate anomaly score for now due to its origin in a multivariate output distribution. [su_robust_2019] propose decomposing the anomaly score into anomaly scores for each channel, by splitting the multivariate output distribution into $d_{\mathcal{D}}$ univariate output distributions. The resulting anomaly scores for each of the $d_{\mathcal{D}}$ channels are referred to as univariate anomaly scores due to their origin in $d_{\mathcal{D}}$ univariate output distributions. To find the channel that contributed most to a detection, the highest univariate anomaly score is assumed to be the root cause.
In this work, the methodology proposed by [su_robust_2019] is used to find the channels that contribute most to a detection. To this end, the method is slightly adapted, given the online premise of the use case. Rather than considering the univariate anomaly scores as time series for root cause analysis, the first time step of the multivariate anomaly score above the threshold is used. Then, for that time step, the highest univariate anomaly score is assumed as the contributing channel.
@@ 5.1 Setup
During training only, input windows are corrupted using Gaussian noise using $0.01$ standard deviation to increase robustness [vincent_extracting_2008].
Key factors that are investigated in Section 5 are given a default value which applies to all experiments unless otherwise specified. These factors are training and validation subset size, which is set to $512$h, reverse-window method, where the mean-type is used, the latent dimension size, which is set to $d_{\textbf{Z}}=64$, the MA mechanism, which is set up as proposed in [vaswani_attention_2017] with a head count of $h=8$ and a key dimension size $d_{\textbf{K}}=\lfloor d_{\mathcal{D}}/h\rfloor=1$.
@@
Cyclical $D_{\text{KL}}$ annealing is applied to the training of TeVAE, to avoid the $D_{\text{KL}}$ vanishing problem [fu_cyclical_2019]. The $D_{\text{KL}}$ vanishing problem occurs when regularisation is too strong at the beginning of training, i.e. the Kullback-Leibler divergence term has a larger magnitude in relation to the reconstruction term. Cyclical $D_{\text{KL}}$ annealing allows the model to weigh the Kullback-Leibler divergence lower than the reconstruction term in a cyclical manner through a weight $\beta$. This callback is configured with a grace period of 25 epochs, where $\beta$ is linearly increased from 00 to $10^{-8}$. After the grace period, $\beta$ is set to $10^{-8}$ and is gradually increased linearly to $10^{-2}$ throughout the following $25$ epochs, representing one loss cycle. This loss cycle is repeated until the training stops.
@@ 5.2 Online Evaluation Metrics
As discussed in Section 1, a testing subset in discrete time-series anomaly detection problem has three types of time series: entirely normal, time-series anomalies and sub-sequence anomalies. Since anomalies are considered rare events, the number of anomaly-free time series $N_{\text{af}}$ within $\mathcal{D}^{\text{test}}$ is much larger than the number of time-series anomalies $N_{\text{ts}}$ and sub-sequence anomalies $N_{\text{ss}}$, such that $N_{\text{af}}>>N_{\text{ts}}+N_{\text{ss}}=N_{\text{a}}$ and $N=N_{\text{af}}+N_{\text{ts}}+N_{\text{ss}}$. In the case of anomaly-free time series and time-series anomalies, traditional labels can easily be applied. An anomaly-free time series can be labelled as true negative or false positive and a time-series anomaly can be labelled as true positive or false negative. For partially anomalous time series, i.e. where the anomalous behaviour occupies a contiguous subset of time steps within the time series, it can be labelled as a true positive or a false negative, but also as a false positive, which occurs when an algorithm flags a time step early. Formally, this is the case when the first flagged time step is far enough ahead of the first ground-truth time step that the model cannot have had access to it. As is evident, the proposed metrics can only be applied to discrete time-series data sets where anomalous time series have at most one contiguous ground-truth anomalous sub-sequence of any length. Also, ensuring each time series contains at most one contiguous anomaly avoids ambiguity on how multiple sub-sequence anomalies within a time series should be detected and counted [wu_current_2021]. In addition to that, there can also be at most one contiguous ground-truth anomaly-free sub-sequence of any length within a sub-sequence anomalous time series, since systems that can dynamically return to anomaly-free behaviour (especially in a short amount of time) reap less benefit from automated anomaly detection than those which require human attention. Given that a predicted anomaly likely requests human attention and potentially a stoppage of the system, the adaptation of the metrics proposed above only takes into account the first predicted anomalous time step. For cases where the process may continue, there can be multiple contiguous predicted anomalous sub-sequences, hence the metrics cannot be applied.
@@
where $K$ is the number of discrete points along the precision-recall curve, $k$ the index of discrete points along the precision-recall curve and $\Delta R_{k}$ the sub-interval length between indices $k$ and $k-1$.

### v2

@@ 4.1 Overview
Figure 3: An illustration of the proposed TeVAE model. Blue shapes designate trainable models, orange deterministic tensors and green distribution parameters. The shape of each tensor is designated below it. During training Z is used as the value matrix, denoted by the solid arrow, whereas during inference $\boldsymbol{\mu}_{\textbf{Z}}$ is used as the value matrix, denoted by the traced arrow. Topologically, TeVAE resembles MA-VAE [7].
@@ 4.2 Inference Mode
Despite the generative capabilities of VAEs, TeVAE does not leverage generation for anomaly detection. Rather than sampling a latent matrix as shown in Equation 14 during inference, sampling is disabled and only $\boldsymbol{\mu}_{\textbf{Z}}$ is taken as the input for the multi-head attention mechanism, like in [22]. Equation 14, therefore, is replaced by Equation 19 for the forward pass.
@@
This not only accelerates inference by eliminating the sampling process but is also empirically found to be a good approximation of an averaged latent matrix if it were sampled several times like in [18]. The TeVAE layout during inference is shown in Figure 3, where the traced arrow designates the information flow from the encoder to the MA mechanism.
@@ 4.4 Bypass Phenomenon
VAE, when combined with an attention mechanism, can exhibit a behaviour called the bypass phenomenon [3]. When the bypass phenomenon happens the latent path between encoder and decoder is bypassed and information flow occurs mostly or exclusively through the attention mechanism, as it has deterministic access to the encoder hidden states and therefore avoids regularisation through the $D_{\text{KL}}$ term. In an attempt to avoid this, [3] propose variational attention, which, like the VAE, maps the input to a distribution rather than a deterministic vector. Applied to natural language processing, [3] demonstrate that this leads to a diversified generated portfolio of sentences, indicating alleviation of the bypassing phenomenon. As previously mentioned, only [18] applies this insight in the anomaly detection domain, however, they do not present any evidence that it alleviates the bypass phenomenon in their work. TeVAE on the other hand, cannot suffer from the bypass phenomenon in the sense that information flow ignores the latent variational path between encoder and decoder since the MA mechanism requires the value matrix V from the encoder to output the context matrix. Assuming the bypass phenomenon also applies to a case where information flow ignores the attention mechanism, one could claim that TeVAE is not immune. To disprove this claim, the attention mechanism is removed from the model in an ablation study to see if anomaly detection performance remains the same. In this case, $\textbf{V}=\textbf{Z}$ is instead directly input into the decoder. If it drops, it is evidence of the contribution of the attention mechanism to the model performance and hence is not bypassed. The results for this ablation study are shown and discussed in Section 5.
@@ 4.5 Reverse-window Process
Since the model is trained to reconstruct fixed-length windows, the same applies during inference. However, to decide whether a given measurement sequence $\mathcal{S}_{n}\in\mathbb{R}^{T_{n}\times d_{\mathcal{D}}}$ is anomalous, a continuous reconstruction of the measurement is required. A trivial way to do so would be to window the input measurement $\mathcal{S}_{n}$ using a shift of $1$, input the windows into the model and chain the last time step from each output window to obtain a continuous sequence [5]. Considering the BiLSTM nature of the encoder and decoder, the first and last time steps of a window can only be computed given the states from one direction, making these values, in theory, less accurate, however. To overcome this, we propose averaging matching time steps in overlapping windows, which is called mean-type reverse-window method. This is done by pre-allocating an array with NaN values, filling it, and taking the mean for each time step while ignoring the NaN values, as depicted in Figure 4.
@@
This process and the general anomaly detection process are described in Algorithm 1. The input for this process is the output distribution parameters $(\boldsymbol{\mu}_{\textbf{X}},\log\boldsymbol{\sigma}_{\textbf{X}}^{2})$, so it essentially averages the distributions and hence can only be applied to the mean and variance parameters. Consider the distributions $\mathcal{N}(\mu_{x},\sigma_{x}^{2})$ and $\mathcal{N}(\mu_{y},\sigma_{y}^{2})$, both assumed to be independant and normally distributed. The sum of both distributions results in normal distribution $\mathcal{N}(\mu_{x+y},\sigma_{x+y}^{2})$, obtained as shown in Equation 20 [13].
@@
$\text{Sequence }\mathcal{S}_{n}\in\mathbb{R}^{T_{n}\times d_{\mathcal{D}}},\text{Threshold }\tau$
$\text{Label }l_{n}\text{ for sequence }\mathcal{S}_{n}$
$n_{\text{windows}}\leftarrow T_{n}-w+1$ $\triangleright$ Find the total number of windows in sequence
$\boldsymbol{\mu}_{\textbf{X},\text{temp}}\leftarrow\text{zeros}(n_{\text{windows}},T_{n},d_{\mathcal{D}})+\text{NaN}$ $\triangleright$ Pre-allocate NaN array to assign all windows to
$\boldsymbol{\sigma}_{\textbf{X},\text{temp}}^{2}\leftarrow\text{zeros}(n_{\text{windows}},T_{n},d_{\mathcal{D}})+\text{NaN}$ $\triangleright$ Pre-allocate NaN array to assign all windows to
@@
  $\boldsymbol{\mu}_{\textbf{X},\text{temp}}[i,i:i+w]\leftarrow\boldsymbol{\mu}_{\textbf{X}}$ $\triangleright$ Assign mean parameter to pre-allocated array
  $\boldsymbol{\sigma}_{\textbf{X},\text{temp}}^{2}[i,i:i+w]\leftarrow\boldsymbol{\sigma}_{\textbf{X}}^{2}$ $\triangleright$ Assign variance parameter to pre-allocated array
@@
$\boldsymbol{\mu}_{\mathcal{S}}\leftarrow\text{nanmean}(\boldsymbol{\mu}_{\textbf{X},\text{temp}},0)$ $\triangleright$ Take nanmean along axis 0
$\boldsymbol{\sigma}_{\mathcal{S}}^{2}\leftarrow\text{nanmean}(\boldsymbol{\sigma}_{\textbf{X},\text{temp}}^{2},0)$ $\triangleright$ Take nanmean along axis 0
@@
The theoretical delay $\delta_{\text{theory}}$ associated with each of the reverse-window processes can be discussed ahead of Section 5, however. $\delta_{\text{theory}}$ is defined as the intrinsic delay introduced by each reverse-windowing method. To illustrate the theoretical delay $\delta_{\text{theory}}$, it is plotted against time $t$ to demonstrate the delay for each of the reverse-window processes, shown in Figure 5. For the last-type reverse-window method during $0<t<w$, no time steps can be evaluated until a full window can be formed, streamed and evaluated, introducing a theoretical delay $\delta_{\text{theory}}=w$. This property is intrinsic to approaches based on fixed-length windows rather than variable-length sequences. At time step $t=w$, however, all time steps $0<t<w$ are evaluated and output at the same time. For $w<t<T$, however, the last time step of each window corresponds the current real-world time step, i.e. $\delta_{\text{theory}}=0$ for $w<t<T$. As mentioned above, the lack of evaluation until $t=w$ is natural to any window-based approach and hence the first-type and mean-type reverse-window methods show the same behaviour. In contrast to last-type reverse-windowing, however, these methods only output the first value of each window i.e. evaluation is $\delta_{\text{theory}}=w$ time steps behind for 0<t<Tw0<t<T-w. At $t=T$, though, the time steps Tw<t<$T$-w<t<T are all output at the same time, meaning that $\delta_{\text{theory}}=0$ and no extra time is needed for evaluation after streaming ends. It should be noted, however, that in online time-series anomaly detection the last-type reverse-window method is in theory faster than the other two types. Consider a sub-sequence anomaly starting at time step $w+2$, designated by the red line in Figure 5. Apart from the time required for inference, the last-type reverse-window method can detect the anomaly without a theoretical delay $\delta_{\text{theory}}$, during $w<t<T$, whereas the other two methods can only detect the anomaly $\delta_{\text{theory}}=w$ time steps later. For 0<t<Tw0<t<T-w, a theoretical delay of $\delta_{\text{theory}}=w$ time steps is, therefore, the absolute best that first-type and mean-type reverse-windowing can achieve, however, while the best delay the first-type can achieve is $\delta_{\text{theory}}=0$, it will be higher in reality, perhaps higher than $w$ time steps.
@@ 4.6 Root-cause Analysis
The first known to the authors is proposed by [28]. The approach does not use time series but instead, the resulting correlation matrices and the anomaly score is given as the difference between the input and reconstructed correlation matrices. Hence, as a mean of providing root-cause information they rank each channel by the number of high anomaly scores for every other channel. Root-cause identification is then quantified by the recall for the top $k$ channels, in this case three, although little information on how exactly this is calculated is provided. Another issue with this metric is that is not parameter-free, since $k$ needs to be set manually. Furthermore, the method used is specific to approaches using the same type of correlation matrices, which is rare in literature.
Approaches using the negative log-likelihood as the anomaly score obtain it from a multivariate output distribution with a diagonal covariance. While the resulting anomaly score is a univariate time series, it is referred to as multivariate anomaly score for now due to its origin in a multivariate output distribution. [24] propose decomposing the anomaly score into anomaly scores for each channel, by splitting the multivariate output distribution into $d_{\mathcal{D}}$ univariate output distributions. The resulting anomaly scores for each of the $d_{\mathcal{D}}$ channels are referred to as univariate anomaly scores due to their origin in $d_{\mathcal{D}}$ univariate output distributions. To find the channel that contributed most to a detection, the highest univariate anomaly score is assumed to be the root cause.
In this work, the methodology proposed by [24] is used to find the channels that contribute most to a detection. To this end, the method is slightly adapted, given the online premise of the use case. Rather than considering the univariate anomaly scores as time series for root cause analysis, the first time step of the multivariate anomaly score above the threshold is used. Then, for that time step, the highest univariate anomaly score is assumed as the contributing channel.
@@ 5.1 Setup
During training only, input windows are corrupted using Gaussian noise using $0.01$ standard deviation to increase robustness [26].
Key factors that are investigated in Section 5 are given a default value which applies to all experiments unless otherwise specified. These factors are training and validation subset size, which is set to 512512h, reverse-window method, where the mean-type is used, the latent dimension size, which is set to $d_{\textbf{Z}}=64$, the MA mechanism, which is set up as proposed in [25] with a head count of $h=8$ and a key dimension size $d_{\textbf{K}}=\lfloor d_{\mathcal{D}}/h\rfloor=1$.
@@
Cyclical $D_{\text{KL}}$ annealing is applied to the training of TeVAE, to avoid the $D_{\text{KL}}$ vanishing problem [9]. The $D_{\text{KL}}$ vanishing problem occurs when regularisation is too strong at the beginning of training, i.e. the Kullback-Leibler divergence term has a larger magnitude in relation to the reconstruction term. Cyclical $D_{\text{KL}}$ annealing allows the model to weigh the Kullback-Leibler divergence lower than the reconstruction term in a cyclical manner through a weight $\beta$. This callback is configured with a grace period of 25 epochs, where $\beta$ is linearly increased from 0 to $10^{-8}$. After the grace period, $\beta$ is set to $10^{-8}$ and is gradually increased linearly to $10^{-2}$ throughout the following $25$ epochs, representing one loss cycle. This loss cycle is repeated until the training stops.
@@ 5.2 Online Evaluation Metrics
As discussed in Section 1, a testing subset in discrete time-series anomaly detection problem has three types of time series: entirely normal, time-series anomalies and sub-sequence anomalies. Since anomalies are considered rare events, the number of anomaly-free time series $N_{\text{af}}$ within $\mathcal{D}^{\text{test}}$ is much larger than the number of time-series anomalies $N_{\text{ts}}$ and sub-sequence anomalies $N_{\text{ss}}$, such that $N_{\text{af}}>>N_{\text{ts}}+N_{\text{ss}}=N_{\text{a}}$ and $N=N_{\text{af}}+N_{\text{ts}}+N_{\text{ss}}$. In the case of anomaly-free time series and time-series anomalies, traditional labels can easily be applied. An anomaly-free time series can be labelled as true negative or false positive and a time-series anomaly can be labelled as true positive or false negative. For partially anomalous time series, i.e. where the anomalous behaviour occupies a contiguous subset of time steps within the time series, it can be labelled as a true positive or a false negative, but also as a false positive, which occurs when an algorithm flags a time step early. Formally, this is the case when the first flagged time step is far enough ahead of the first ground-truth time step that the model cannot have had access to it. As is evident, the proposed metrics can only be applied to discrete time-series data sets where anomalous time series have at most one contiguous ground-truth anomalous sub-sequence of any length. Also, ensuring each time series contains at most one contiguous anomaly avoids ambiguity on how multiple sub-sequence anomalies within a time series should be detected and counted [27]. In addition to that, there can also be at most one contiguous ground-truth anomaly-free sub-sequence of any length within a sub-sequence anomalous time series, since systems that can dynamically return to anomaly-free behaviour (especially in a short amount of time) reap less benefit from automated anomaly detection than those which require human attention. Given that a predicted anomaly likely requests human attention and potentially a stoppage of the system, the adaptation of the metrics proposed above only takes into account the first predicted anomalous time step. For cases where the process may continue, there can be multiple contiguous predicted anomalous sub-sequences, hence the metrics cannot be applied.
@@
where $K$ is the number of discrete points along the precision-recall curve, $k$ the index of discrete points along the precision-recall curve and $\Delta R_{k}$ the sub-interval length between indices $k$ and k1k-1.